### v2

# GRAPES: Learning to Sample Graphs for Scalable Graph Neural Networks

Abstract
Graph neural networks (GNNs) learn to represent nodes by aggregating information from their neighbors. As GNNs increase in depth, their receptive field grows exponentially, leading to high memory costs. Several existing methods address this by sampling a small subset of nodes, scaling GNNs to much larger graphs. These methods are primarily evaluated on homophilous graphs, where neighboring nodes often share the same label. However, most of these methods rely on static heuristics that may not generalize across different graphs or tasks. We argue that the sampling method should be adaptive, adjusting to the complex structural properties of each graph. To this end, we introduce GRAPES, an adaptive sampling method that learns to identify the set of nodes crucial for training a GNN. GRAPES trains a second GNN to predict node sampling probabilities by optimizing the downstream task objective. We evaluate GRAPES on various node classification benchmarks, involving homophilous as well as heterophilous graphs. We demonstrate GRAPES’ effectiveness in accuracy and scalability, particularly in multi-label heterophilous graphs. Unlike other sampling methods, GRAPES maintains high accuracy even with smaller sample sizes and, therefore, can scale to massive graphs. Our code is publicly available at https://github.com/dfdazac/grapes.
1 Introduction
In many applications, data is represented with graph structures, such as in recommender systems, social networks, and the chemical and medical domains [31, 42, 25]. In these domains, graph neural networks (GNNs) are a powerful tool for representation learning on graphs [23, 38, 46].
Despite the broad range of applications of GNNs, scalability remains a significant challenge [35]. Unlike traditional machine learning problems where data is assumed to be i.i.d., the graph structure introduces dependencies between a node and its neighborhood. This complicates partitioning data into mini-batches. Additionally, the number of nodes that a GNN needs to process increases exponentially with the number of layers.
Most current graph sampling methods use a fixed heuristic to compute node inclusion probabilities: The sampling process is independent of the node features, graph structure, or the GNN task performance [9, 51, 47]. For simple graphs like homophilous graphs [50, 49, 48] where there is a strong correlation between labels of a node and its neighbors, a fixed heuristic typically suffices. However, these methods fall short when applied to graphs in which only a few of the nodes in the neighborhood provide relevant information for the task at hand. Consequently, we study adaptive sampling, where the sampling method adapts dynamically to the task by learning which nodes should be included. Our adaptive sampling technique learns to sample by directly minimizing the classification loss of the downstream GNN. This differs from most existing graph sampling methods, which aim to approximate the full-batch GNN behavior [9, 51, 47, 18], opting for an indirect approach to achieve high accuracy.
Figure 1: Overview of GRAPES. First, GRAPES processes a target node (green) by computing node inclusion probabilities on its 1-hop neighbors (shown by node color shade) with a sampling GNN. Given these probabilities, GRAPES samples $k$ nodes. Then, GRAPES repeats this process over nodes in the 2-hop neighborhood. We pass the sampled subgraph to the classifier GNN for target node classification. Finally, GRAPES uses the classification loss to update the classifier GNN and to reward the sampler GNN.
We introduce a straightforward adaptive sampling method called Graph Adaptive Sampling (GRAPES). As illustrated in Figure 1, GRAPES samples a subgraph around the target nodes in a series of steps. At each step, a sampling policy GNN computes inclusion probabilities for the nodes neighboring the current subgraph, which the sampler uses to select a subset. Once the sampling is complete, the resulting subgraph is passed to a second classifier GNN for classification. The classification loss is then backpropagated to train both GNNs.
To enable backpropagation through the sampling step, we need a gradient estimator. We compare a reinforcement learning (RL) approach and a GFlowNet (GFN) approach [6]. Both approaches allow us to consider the sampling process and the GNN computation together and to train them concurrently. This allows the sampler to adapt based on factors like the node features, the graph structure, the sample size, and other contextual features.
In addition to the datasets commonly used in the literature, we evaluate GRAPES on several heterophilous and/or multi-label graphs and demonstrate its effectiveness on these complicated graphs. Moreover, we spend significant effort on the evaluation protocol for fairly testing graph sampling methods. In particular, we ensure that all sampling methods are evaluated under the same conditions and on the same GNN architecture to eliminate any confounding factors and to ensure that any changes in performance can only be attributed to the sampling method. To the best of our knowledge, we are the first to perform such a rigorous comparison on twelve varied datasets.
We evaluate GRAPES on several node classification tasks and find that
  1. 1.
    Competitive Performance. GRAPES is competitive to other sampling methods in accuracy across datasets. On heterophilous and multi-label datasets GRAPES provides state-of-the-art performance.
  2. 2.
    Homophilous Graphs. For some benchmarks, mostly those based on homophilous graphs, simple baselines like random sampling perform as well as any state-of-the-art method, including GRAPES.
  3. 3.
    Memory Efficiency.  GRAPES is competitive with a state-of-the-art non-sampling GNN scaling method, while using up to an order of magnitude less GPU memory.
2 Related Work
Fixed Sampling Policy: In this category, the sampling policy is independent of the training of the GNN and is based on a fixed heuristic that does not involve any training. Given a set of target nodes, i.e. the nodes that are to be classified, node-wise sampling methods sample a given number of nodes for each target node. GraphSage [16] is a node-wise sampling method that randomly samples nodes. However, node-wise sampling can result in nodes being sampled multiple times redundantly because it can be the neighbor of several nodes [51]. A more efficient approach is layer-wise sampling, for example, FastGCN [9] and LADIES [51]. They aim to minimize variance by sampling nodes in each layer with probabilities proportional to their degree. Moreover, some techniques focus on sampling subgraphs in each mini-batch, like GraphSAINT [47] and ClusterGCN [10]. While these techniques effectively scale GNNs to larger graphs, they do not adapt to the sampling policy based on the GNN’s performance on the task. In graph signal processing community, the authors in [15] propose a node sampling technique based on [4] that aims to sample nodes for a unique and stable graph signal reconstruction. However, this method is only applied to small graphs.
Learnable Sampling Policy: A few methods learn the probability of including a node based on feedback from the GNN. AS-GCN [18] is a method that learns a linear function that estimates the node probabilities layer-wise. Similarly, PASS [44] learns a mixture of a random distribution and a learned policy with RL. Like GRAPES, it is task-adaptive, however, its sampling probability uses a bilinear similarity measure between neighboring node features, meaning the graph structure is not considered. GNN-BS [27] formulates the node-wise sampling problem as a bandit problem and updates the sampling policy according to a reward function that reduces the sampling variance. SubMix [1] proposes a mixture distribution of sampling heuristics with learnable mixture weights. DSKReG [40] learns the relevance of items in a user-item knowledge graph by jointly optimizing the sampling strategy and the recommender model. The majority of these methods focus on variance reduction and fail to consider the classification loss, unlike GRAPES. We argue that adaptivity to the classification loss, allows for sampling the influential nodes depending on the task, and results in better performance.
Other Scalable Methods: Authors of [33] propose a method to transfer the weights of a GNN trained on a mid-sized graph to larger graphs given the graphon similarity between the graphs. Another group of papers uses historical embeddings of the nodes when updating the target nodes’ embeddings [8, 13, 45, 37]. GAS [13] approximates the embeddings of the 1-hop neighbors using the historical embeddings of those nodes learned in the previous training iterations. These methods reduce the GPU memory usage by training in mini-batches and learning from the 1-hops neighbors with the historical embeddings saved in CPU memory. Unlike GRAPES, they process the whole graph.
3 Background: GNN Training and Sampling
We first provide the necessary background about GNNs and graph sampling. Although our method is independent of the choice of GNN architecture, we limit our discussion to the GCN architecture for simplicity [23].
Let $\mathcal{G}=(\mathcal{V,E})$ be an undirected graph with a list of $N$ nodes $\mathcal{V}=\{1,\ldots,N\}$ and a set of edges $\mathcal{E}$. The adjacency matrix $A\in\{0,1\}^{N\times N}$ indicates a connection between a pair of nodes. Next, let $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ and where $\tilde{D}$ is the degree matrix of $\tilde{A}$. Let $X\in\mathbb{R}^{N\times f}$ be the node embeddings and let $Y$ be the labels for the target nodes $\mathcal{V}^{t}\subset\mathcal{V}$, where $\mathcal{V}^{t}$ indexes the nodes with a label. In node classification, we want to learn to predict the labels of the target nodes.
We consider a GCN with $L$ layers. The output of the $l$-th layer of the GCN is $H^{(l)}=\sigma(\tilde{A}H^{(l-1)}W^{(l)})$, where $W^{(l)}$ is the weight matrix of GCN layer $l$ and $\sigma$ is a non-linear activation function. For a node $i\in\mathcal{V}$, this corresponds to the update
$h^{(l)}_{i}=\sigma\left(\sum_{j\in\mathcal{N}(i)\cup\{i\}}\hat{A}_{ij}h^{(l-1)}_{j}W^{(l)}\right),$ (1)
where $\mathcal{{N}}(i):\mathcal{V}\rightarrow 2^{\mathcal{V}}$ is the set of $i$’s neighbors excluding $i$.
As the number of layers increases, the computation of the embedding of the node $i$ involves neighbors from further hops. As a result, the neighborhood size grows rapidly with the number of layers. We study how to sample the graph to overcome this growth. We focus on layer-wise sampling, which is a common type of graph sampling approaches [9, 51, 18]. First, we divide the target nodes in mini-batches of size $b$. Then, in each layer, we sample $k$ nodes $\mathcal{V}^{(l)}$ among the neighbors of the nodes in the previous layer using the sampling policy $q$. To make this precise, we will need some additional notation. The computation of the output of layer $l$ is then:
for all $i\in K^{(l)}$, $\quad h^{\prime(l)}_{i}=\sigma\left(\sum_{j\in K^{(l-1)}}\hat{A}^{(l)}_{ij}h^{\prime(l-1)}_{j}W^{(l)}\right)$ (2)
with $\quad\mathcal{V}^{(l-1)},\mathcal{V}^{(l)}\sim q(\mathcal{V}^{(l-1)},\mathcal{V}^{(l)}|k),$
where
  1. 1.
    $K^{(0)}=\mathcal{V}^{(0)}$ is the set of target nodes in the current mini-batch (a subset of $\mathcal{V}^{t}$);
  2. 2.
    $K^{(l)}=\mathcal{V}^{(l)}\cup\mathcal{V}^{(0)}$ for all $l\in 1,\ldots,L$ adds the batch nodes $\mathcal{V}^{(0)}$ to the sampled nodes $\mathcal{V}^{(l)}$ to ensure self-loops between the batch nodes;
  3. 3.
    $\mathcal{V}^{(l)}\subseteq\mathcal{N}(K^{(l-1)})$ are the nodes sampled in layer $l$ among the neighbors $\mathcal{N}(K^{(l-1)})$ of the nodes in $K^{(l-1)}$. Note that $\mathcal{V}^{(l)}$ cannot contain nodes in $K^{(l-1)}$;
  4. 4.
    $\hat{A}^{(l)}=D^{(l)^{-1/2}}A^{(l)}D^{(l)^{-1/2}}$ is computed from the adjacency matrix $A^{(l)}$ containing the edges between $K^{(l)}$ and $K^{(l-1)}$, and the corresponding degree matrix $D^{(l)}$. [^1] Unlike the full-batch GCN, the adjacency matrix varies across layers when sampling because each layer involves a different set of nodes. Note also that, unlike the full-batch setting, message passing is asymmetric: node $i$ may be updated from node $j$ but not vice versa.To be precise, the entries $A^{(l)}_{ij}$ are 1 if and only if $i\in K^{(l)}$, $j\in K^{(l-1)}$ and if there is an edge $(i,j)$ in the original graph, that is, $\tilde{A}_{ij}=1$;
  5. 5.
    $q$ is a sampling policy that samples the $k$ nodes $\mathcal{V}^{(l)}$.
Existing layer-wise sampling methods such as LADIES or FastGCN use a fixed heuristic to determine $q$, for instance, by computing node probabilities proportional to the node degrees. However, an adaptive method learns the distribution $q$ instead.
4 Graph Adaptive Neighbor Sampling (GRAPES)
In this section, we introduce Graph Adaptive Sampling (GRAPES). GRAPES is a layer-wise and layer-dependent sampling method. In each layer $l$, we sample a subset $\mathcal{V}^{(l)}$ that is much smaller than the neighborhood of the previous layer. That is, $|\mathcal{V}^{(l)}|=k\ll|\mathcal{N}(K^{(l-1)})|$, where again $K^{(l-1)}=\mathcal{V}^{(0)}\cup\mathcal{V}^{(l-1)}$ adds the batch nodes to the sampled nodes of the previous layer. We use a second GNN to compute the inclusion probability for each node in $\mathcal{N}(K^{(l-1)})$. In the remainder of this section, we describe our sampling policy $q$ and training methods for the sampling policy. Furthermore, Algorithm 1 shows one epoch of GRAPES in pseudocode.
4.1 Sampling policy
Next, we specify the sampling policy GNN $\operatorname{GCN}_{\operatorname{S}}(K^{(l-1)})$, which computes inclusion probabilities on the subgraph created from the nodes in $K^{(l-1)}$ and their neighbors $\mathcal{N}^{(l-1)}$. The sampling policy $q$ decomposes as $q(\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}|\mathcal{V}^{(0)})=\prod_{l=1}^{L}q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$. We compute each factor $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ as a product of Bernoulli inclusion probabilities, given by $\operatorname{GCN}_{\operatorname{S}}$, for each node $i$ in the neighborhood $\mathcal{N}(K^{(l-1)})$:
$q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})=\prod_{i\in\mathcal{N}(K^{(l-1)})}\operatorname{Bern}(i\in\mathcal{V}^{(l)}|p_{i}),\quad p_{i}=\operatorname{GCN}_{\operatorname{S}}(K^{(l-1)})_{i}.$ (3)
In addition to the regular embeddings $X$, the sampler GNN also has access to a one-hot vector of length $L+1$ that records the value $l+1$ for nodes sampled in layer $l$ (with 1 recorded for the target nodes). This allows the sampler to differentiate between nodes sampled in different layers.
4.2 Sampling exactly $k$ nodes
We have a clear constraint on the number of nodes we want to include in training: In Equation 2, we sample exactly $k$ nodes without replacement. However, our sampling policy $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ consists of many independent Bernoulli distributions, and it is highly unlikely that we sample exactly $k$ nodes from this distribution.
Instead, we use the Gumbel-Top-$k$ trick [39, 19], which selects a set of exactly $k$ nodes $\mathcal{V}^{(l)}$ by perturbing the log probabilities randomly and taking the top-$k$ among those:
$\mathcal{V}^{(l)}=\operatorname*{\operatorname{top-\!}k}_{i\,\in\,\mathcal{N}(K^{(l-1)})}\log p_{i}+\epsilon_{i},\quad\epsilon_{i}\sim\operatorname{Gumbel}(0,1)$ (4)
This guarantees a sample from $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$ that conditions on the number of nodes sampled.
While this results in a tractable and adaptive sampling procedure of exactly $k$ nodes for the classifier GNN, we have not yet given a method for learning the sampling policy. Unfortunately, a sampling operation, or in this case, the top-$k$ operation, provides no functional gradient [30]. We will resort to simple methods from the reinforcement learning and GFlowNet literature to still be able to train the sampling policy, which we will explain next.
4.3 Training the Sampling Policy
We train the sampling policy to minimize the classification loss $\mathcal{L}_{\operatorname{C}}$ of the classifier GNN [^2] GRAPES can be extended to other tasks, but we focus on node classification in the current work.. We use two methods to train the sampling policy $\operatorname{GCN}_{\operatorname{S}}$: a reinforcement learning (RL) method and a GFlowNet (GFN) method.
REINFORCE (GRAPES-RL)
In the REINFORCE-based method, we use a simple REINFORCE estimator [41] to compute an unbiased gradient of the classification loss.
$\mathcal{L}_{\operatorname{RL}}(X,Y,\mathcal{V}^{(0)})=\mathcal{L}_{\operatorname{C}}(X,Y,K^{(0)},\ldots,K^{(L)})\log q(\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}|\mathcal{V}^{(0)}),$ (5)
where we sample $\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}\sim q(\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}|\mathcal{V}^{(0)},k)$. Taking the derivative with respect to the parameters of $q$ results in the standard REINFORCE estimator. Note that this is an off-policy estimator since we sample from $q$ conditioned on the number of samples $k$ with the Gumbel-Top-$k$ trick, but compute gradients with respect to the distribution unconditioned on $k$. This is because computing likelihoods conditioned on $k$, although possible [3], is computationally expensive. We discuss this issue in more detail in Appendix A.
GFlowNets (GRAPES-GFN)
The second method uses the Trajectory Balance loss [28] from the GFlowNet literature [7, 6], which is known to perform well in off-policy settings [29].
$\mathcal{L}_{\operatorname{GFN}}(X,Y,\mathcal{V}^{(0)})=\Big{(}$ $\log Z(\mathcal{V}^{(0)})+\log q(\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(L)}|\mathcal{V}^{(0)})$ (6)
$+\alpha\cdot\mathcal{L}_{\operatorname{C}}(X,Y,K^{(0)},\ldots,K^{(L)})\Big{)}^{2},$
where $\log Z(\mathcal{V}^{(0)})$ is a small GCN that predicts a scalar from the target nodes, and $\alpha$ is a tunable reward scaling hyperparameter. For a detailed derivation, see Appendix F.3. GFlowNets minimize an objective that ensures sampling in proportion to the negative classification loss, rather than minimizing it like in REINFORCE. This may have benefits in the exploratory behavior of the sampler, as it encourages the training of diverse sets of nodes, instead of only the single best set of nodes.
Algorithm 1 One GRAPES epoch
1:Graph $\mathcal{G}$, node embeddings $X$, node labels $Y$, target nodes $\mathcal{V}^{t}$, batch size $b$, sample size $k$, GCN for classification $\operatorname{GCN}_{\operatorname{C}}$, and GCN for sampling policy $\operatorname{GCN}_{\operatorname{S}}$.
2:Divide target nodes $\mathcal{V}^{t}$ into batches $\mathcal{V}^{(0)}$ of size $b$
3:for each batch $\mathcal{V}^{(0)}$ do
4:     $K^{(0)}\leftarrow\mathcal{V}^{(0)}$
5:     for layer $l=1$ to $L$ do
6:         for node $i$ in $\mathcal{N}(K^{(l-1)})$ do
7:              $p_{i}\leftarrow\operatorname{GCN}_{S}(K^{(l-1)})_{i}$triangleright Compute probabilities of inclusion
8:              $\epsilon_{i}\sim\operatorname{Gumbel}(0,1)$triangleright Sample Gumbel noise
9:         $\mathcal{V}^{(l)}\leftarrow\operatorname*{\operatorname{top-\!}k}_{i\in\mathcal{N}(K^{(l-1)})}\log p_{i}+\epsilon_{i}$triangleright Get $k$ best nodes (Eq. 4)
10:         $K^{(l)}\leftarrow\mathcal{V}^{(0)}\cup\mathcal{V}^{(l)}$triangleright Add target nodes
11:     $\ell_{\operatorname{C}}\leftarrow\mathcal{L}_{\operatorname{C}}(X,Y,K^{(0)},\ldots,K^{(L)})$triangleright Compute classification loss
12:     Compute sampling policy loss $\ell_{\operatorname{S}}$ from Eq. 5 or 6
13:     Update parameters of $\operatorname{GCN}_{\operatorname{S}}$ by minimizing $\ell_{\operatorname{S}}$
14:     Update parameters of $\operatorname{GCN}_{\operatorname{C}}$ by minimizing $\ell_{\operatorname{C}}$
5 Experiments
Our experiments aim at answering the following research question: given a fixed sampling budget and GNN architecture, what is the effect of training with an adaptive policy for layer-wise sampling, in comparison with the related work? While several works in the literature of sampling for GNNs have focused on classification benchmarks to demonstrate the performance of sampling algorithms, there are several confounding factors in their experimental setup that prevent proper understanding whether the perceived performance improvements are a result of the sampling method itself. Examples of such confounding factors found in related work include using different architectures, like the GCN [23] or GAT [38], different sizes for the hidden layers, number of layers, batch sizes, number of nodes sampled per layer, and the number of training epochs. These are factors independent of sampling algorithms that nonetheless affect the performance in the benchmarks. We present a detailed overview of differences in the experimental setup of the related work in Table 3 in the appendix.
To better understand the effect of sampling algorithms, we thus carry out experiments on a fixed GCN architecture [23] with two layers, a hidden size of 256, batch size of 256, sample size of 256 nodes per layer, and a fixed number of epochs per dataset (detailed in Appendix B). Under this setting, we compare GRAPES with the following baselines: a Random baseline that uses the same setup as GRAPES but with uniform inclusion probabilities; FastGCN [9], LADIES [51], GraphSAINT [47], GAS [13], and AS-GCN [18]. With this setup, we aim to control our experiments in a way such that variations in performance can only be attributed to the sampling method.
For all baselines (except Random) we rely on their publicly available implementations. For all methods, we optimize the learning rate using the performance on the validation set. GRAPES requires selecting two additional hyperparameters: the learning rate for the sampling policy and, for GRAPES-GFN, the reward scaling parameter $\alpha$. We tune these based on the classification performance on the validation set. These hyperparameters are specific to our sampling method and do not explicitly affect the learning capacity of the GCN classifier.
We refer the reader to Appendix B for more details on hyperparameter settings.
5.1 Datasets
Most methods in the literature of sampling in GNNs are evaluated on homophilous graphs (where similar nodes are likely to be connected by an edge) and multi-class classification, where nodes are classified into one of several classes (in contrast to multi-label). To further understand the effect of sampling, we carry out experiments with heterophilous graphs, where nodes that are connected differ in their features and labels, and multi-label datasets in which a node can be assigned one or more labels. As shown in the literature [50, 48], we argue that node classification with GCN is more challenging on these datasets compared to multi-class classification on homophilous graphs, where an adaptive policy could be beneficial. In particular, we run experiments with the following datasets for the node classification task:
  • Homophilous graphs: citation networks (Cora, Citeseer, Pubmed with the “full” split) [34, 43], Reddit [16], ogbn-arxiv and ogbn-products [17], and DBLP [48].
  • Heterophilous graphs: Flickr [47], Yelp [47], ogbn-proteins [17], BlogCat [48], and snap-patents [24] .
We included the label homophily ratio of all the datasets in Table 1 and 2. For the multi-label datasets, we followed the same framework in [48] to calculate their label homophily ratio. Statistics about the datasets can be found in Appendix C.
Table 1: F1-scores (%) for different sampling methods trained on homophilous graphs, for a batch size of 256, and sample size of 256 per layer. We report the mean and standard deviation over 10 runs. The best values among the sampling baselines (all except GAS) are in bold, and the second best are underlined. MC stands for multi-class and ML stands for multi-label classification.
Dataset Cora Citeseer Pubmed Reddit ogbn-arxiv ogbn-products DBLP
Homophily $h=0.81$ $h=0.74$ $h=0.80$ $h=0.78$ $h=0.65$ $h=0.81$ $h=0.76$
Task MC MC MC MC MC MC ML
GAS 87.00 $\pm$ 0.19 85.87 $\pm$ 0.19 87.45 $\pm$ 0.23 94.75 $\pm$ 0.04 68.36 $\pm$ 0.55 74.69 $\pm$ 0.14 83.08 $\pm$ 0.31
FastGCN 76.17 $\pm$ 3.98 62.81 $\pm$ 7.19 53.52 $\pm$ 28.48 62.93 $\pm$ 3.28 39.49 $\pm$ 8.04 66.09 $\pm$ 3.04 62.93 $\pm$ 3.28
LADIES 76.02 $\pm$ 11.69 63.48 $\pm$ 12.21 72.81 $\pm$ 17.67 59.30 $\pm$ 2.69 43.52 $\pm$ 8.03 68.08 $\pm$ 1.95 59.97 $\pm$ 10.45
GraphSAINT 87.28 $\pm$ 0.49 77.28 $\pm$ 0.67 87.45 $\pm$ 0.75 91.47 $\pm$ 0.94 63.54 $\pm$ 1.75 67.66 $\pm$ 0.69 77.09 $\pm$ 0.56
AS-GCN 85.60 $\pm$ 0.54 79.21 $\pm$ 0.19 90.58 $\pm$ 0.40 93.52 $\pm$ 0.40 65.38 $\pm$ 1.80 73.94 $\pm$ 0.40 83.24 $\pm$ 0.51
Random 86.58 $\pm$ 0.33 78.29 $\pm$ 0.52 90.09 $\pm$ 0.17 94.16 $\pm$ 0.06 61.35 $\pm$ 0.32 70.47 $\pm$ 0.32 76.87 $\pm$ 0.24
GRAPES-RL (ours) 87.62 $\pm$ 0.48 78.75 $\pm$ 0.35 89.40 $\pm$ 0.34 94.09 $\pm$ 0.05 62.58 $\pm$ 0.64 71.45 $\pm$ 0.20 76.88 $\pm$ 0.33
GRAPES-GFN (ours) 87.29 $\pm$ 0.32 78.75 $\pm$ 0.33 89.46 $\pm$ 0.34 94.30 $\pm$ 0.06 61.86 $\pm$ 0.51 70.66 $\pm$ 0.30 77.14 $\pm$ 0.48
Table 2: F1-scores (%) for different sampling methods trained on heterophilous graphs for a batch size of 256, and a sample size of 256 per layer. We report the mean and standard deviation over 10 runs. The best values among the sampling baselines (all except GAS) are in bold, and the second best are underlined. MC stands for multi-class and ML stands for multi-label classification.
Dataset Flickr snap-patents Yelp ogbn-proteins BlogCat
Homophily $h=0.31$ $h=0.22$ $h=0.22$ $h=0.15$ $h=0.1$
Task MC MC ML ML ML
GAS 49.96 $\pm$ 0.28 38.04 $\pm$ 0.20 37.81 $\pm$ 0.07 7.55 $\pm$ 0.01 6.07 $\pm$ 0.04
FastGCN 46.40 $\pm$ 2.51 29.68 $\pm$ 0.61 29.29 $\pm$ 4.39 7.80 $\pm$ 0.55 6.44 $\pm$ 0.41
LADIES 47.19 $\pm$ 3.42 29.09 $\pm$ 1.88 18.92 $\pm$ 3.18 4.31 $\pm$ 0.11 6.74 $\pm$ 0.67
GraphSAINT 48.01 $\pm$ 1.44 28.01 $\pm$ 0.57 34.68 $\pm$ 0.70 9.94 $\pm$ 0.07 6.89 $\pm$ 0.96
AS-GCN 48.42 $\pm$ 1.20 31.04 $\pm$ 0.19 38.51 $\pm$ 1.45 5.20 $\pm$ 0.36 5.43 $\pm$ 0.54
Random 49.39 $\pm$ 0.23 29.74 $\pm$ 0.34 40.63 $\pm$ 0.13 10.82 $\pm$ 0.05 7.13 $\pm$ 0.97
GRAPES-RL (ours) 49.54 $\pm$ 0.67 26.26 $\pm$ 0.50 40.69 $\pm$ 0.55 11.78 $\pm$ 0.14 9.06 $\pm$ 0.68
GRAPES-GFN (ours) 49.29 $\pm$ 0.32 28.46 $\pm$ 0.69 44.57 $\pm$ 0.88 11.57 $\pm$ 0.18 9.22 $\pm$ 0.40
5.2 Results
Comparison with sampling methods We present the F1 scores of GCNs trained via sampling for GRAPES and the sampling-based baselines in Table 1 and 2. We observe that on the majority of heterophilous datasets, and all the heterophilous multi-label datasets, both variations of GRAPES achieve the highest F1-score. As we mentioned earlier, node classification on heterophilous graphs is a challenging task for GCNs, due to the diversity of the neighbor nodes’ labels. Our results suggest that GRAPES’ ability to adapt the sampling policy to particular features of the data allows GRAPES to learn the complex patterns across the heterophilous datasets. Table 1 shows that on most of the homophilous graphs, AS-GCN achieves the highest F1-score, indicating the importance of adaptive sampling. However, unlike GRAPES, this method is only adaptive to the sampling variance and fails to outperform GRAPES on the heterophilous graphs.
FastGCN and LADIES, which use a fixed policy, fail to compete with the other methods in our experimental setup. They assign a higher probability to nodes with higher degrees. This heuristic results in neglecting informative low-degree nodes. Although outperforming LADIES and FastGCN, GraphSAINT, which also uses a fixed sampling policy, fails to outperform the more competitive baselines.
We observe that on some datasets, Random sampling achieves a comparable F1-score as GRAPES and the baselines. This mainly happens on simple datasets, such as homophilous graphs, indicting that on these graphs a simple sampling approach is sufficient.
Comparison with a non-sampling scalable method GAS [13] is a non-sampling method that uses the historical embeddings of the 1-hop neighbors of the target nodes. We present classification results in Table 2, and we visualize memory usage in Fig. 2. While GAS has higher classification F1-scores for some datasets, GRAPES achieves comparable F1-scores and significantly outperforms GAS for Yelp, ogbn-proteins, and BlogCat, all heterophilous multi-label graphs. Once again, these results indicate the effectiveness of the adaptivity of GRAPES in sampling influential nodes in complex graphs. For the memory usage, we compare GAS with the GFlowNet version of GRAPES-32 and GRAPES-256, where $32$ and $256$ (the default value) indicate the sample size per layer. The memory usage results show that, even with a large sample size, GRAPES can use up to an order of magnitude less GPU memory than GAS, especially for large datasets. In large, densely connected graphs such as Reddit, the 1-hop neighborhood can be massive. Then, the difference in memory use for GRAPES, which only sees a small set of neighbors, and GAS, which uses all the neighbors, is significant. While GAS occasionally achieves higher F1-scores, it consistently demands more memory, indicating a potential compromise between accuracy and computational efficiency. In contrast, GRAPES strikes a balance, delivering comparable F1-score with more modest memory footprints.
Figure 2: GPU peak memory allocation (MB) for GAS, and GRAPES-GFN-32 and GRAPES-GFN-256.
Figure 3: Comparative analysis of classification accuracy across different sampling sizes for sampling baseline and GRAPES. We repeated each experiment five times: The shaded regions show the 95% confidence intervals.
GRAPES is robust to low sample sizes. A desirable property of sampling methods, in contrast with full-batch GNNs or methods like GAS (which relies on historical embeddings), is the ability to control the sample size to reduce memory usage as needed. To study this property, we show in Figure 3 the effects of varying the sample size on Reddit, Flickr, Yelp, and ogbn-arxiv. Our results show that both the RL and GFN variants of GRAPES are robust to low sample size, and achieve strong performance with fewer sampled nodes needed than the baselines, enabling training GCNs on larger graphs while using less GPU memory. Random sampling also exhibits robustness to sample size for all datasets except Yelp, where accuracy drops in larger sample sizes. GRAPES-RL shows the same behavior while performing slightly better than Random on Yelp. AS-GCN and GraphSAINT show the largest dependence on sample size, especially on Flickr and Yelp.
GRAPES learns strong preferences over nodes. The ability to selectively choose influential nodes is a crucial property of GRAPES. Figure 4 shows the mean and standard deviation of base 2 entropy for the node preference probabilities for the two layers of $\operatorname{GCN}_{\operatorname{S}}$ for ogbn-products and DBLP for GRAPES-GFN. The probabilities show preference towards particular nodes with a Bernoulli distribution. A well-trained model must have a high preference (probability close to 1) for some nodes and a low preference (probability close to 0) for the rest. Therefore, we would like a low average entropy with a high standard deviation. As the figure shows, the mean entropy in both layers decreases from almost 1 and converges to a value above 0 while the standard deviation increases. This indicates that for ogbn-products, the sampling policy initially assigned a probability near $0.5$, indicating little preference. However, after several training epochs, GRAPES starts preferring some nodes, resulting in lower mean entropy. We observe similar behavior for most datasets (DBLP, BlogCat, Yelp, ogbn-arxiv, and ogbn-proteins). However, we observe that the sampling policy exhibits no preferences among the nodes for the other datasets. For more details about the other datasets, see Appendix D.
Figure 4: Entropy for the ogbn-products and DBLP datasets. The mean is the entropy in bits of the node probability, averaged over nodes. The shaded region indicates the standard deviation of the entropy over all nodes.
6 Discussion and Conclusion
We propose GRAPES, an adaptive graph sampling method based on reinforcement learning and GFlowNet, facilitating the scalability of training GNNs on massive graphs. GRAPES learns the importance of the nodes by adapting to the classification loss. We show how a sampling policy can build a subgraph of influential nodes by learning node preferences that adapt to node features, GNN architecture, classification task, and graph topology. Our experiments demonstrate that GRAPES effectively selects nodes from large-scale graphs and achieves consistent performance over state-of-the-art sampling methods, especially on complex graph structures such as heterophilous graphs. Compared to the other sampling methods, GRAPES can maintain high classification accuracy even with lower sampling rates, indicating GRAPES’ ability to scale to larger graphs by sampling a small but influential set of nodes. GRAPES achieves comparable performance to GAS, while using up to an order of magnitude less memory.
Lack of Uniform Evaluation Protocol. Existing methods in the literature report performance on graph sampling under settings with different GCN architectures, regularization techniques, feature normalization strategies, and data splits, among other differences. These differences made it challenging to determine the benefits of each sampling method. This motivated us to implement a unified protocol across all methods, where we keep the architecture fixed. We encourage future work to consider a similar methodology for a fair evaluation, or an experimental review study, as is common in other areas of machine learning research on graphs [36, 32].
Known Limitations. In our experiments, we only focused on the problem of node classification. However, GRAPES is not tied to a particular downstream task. GRAPES assumes access to a tractable reward function [7]. Therefore, GRAPES will be applicable for other graph-related tasks, like link prediction and unsupervised representation learning. Another limitation is that we only evaluate GRAPES on a GCN architecture with a fixed number of layers. Additionally, there is room for a more thorough investigation into the properties of the subgraphs that GRAPES samples. The experiments at the end of Section 5.2 are a first step towards this. We leave these directions for future work.
Acknowledgments
This project was partially funded by Elsevier’s Discovery Lab and Huawei DREAMS Lab. Emile van Krieken was funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1). We thank Michael Cochez and Ruud van Bakel for insightful discussions. We also thank the Das6 cluster for computing support [5].
Appendix A Off-Policy Sampling Setup
In this Appendix, we discuss the technical and mathematical challenges around our setup that resulted in our off-policy learning setup. In each layer $l$ of the GFlowNet, we aim to sample exactly $k$ out of $n$ nodes. An initially natural setup would be to use the distribution over $k$-subsets of $\mathcal{N}(K^{(l-1)})$ [3]. Using Bayes theorem,
$q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)=\frac{I[|\mathcal{V}^{(l)}|=k]q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})}{\sum_{\mathcal{V}^{{}^{\prime}(l+1)}}I[|\mathcal{V}^{{}^{\prime}(l+1)}|=k]q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})}.$ (7)
When conditioned on $k$, $q$ assigns 0 probability to sets of nodes $\mathcal{V}^{(l)}$ that do not sample exactly $k$ new nodes (that is, when $|\mathcal{V}^{(l)}|\neq k$). However, this requires renormalizing the distribution, which is the function of the denominator term on the right-hand side. Note that this sum is over an exponential number of elements, namely $2^{|\mathcal{N}(K^{(l-1)})|}$, and naive computation is clearly intractable. SIMPLE [3] provides an optimized dynamic programming algorithm for computing this normalization constant. However, it scales polynomially in $|\mathcal{N}(K^{(l-1)})|$ and $k$, and in our experiments, computing the normalizer is a bottleneck already for mid-sized graphs like Reddit.
Therefore, we decided to circumvent having to compute $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$ by sampling using the Gumbel-Top-k trick (Equation 4) to ensure we always add exactly $k$ nodes. However, we are now in an off-policy setting: The samples using Equation 4 are distributed by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$, not by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$, and so we sample from a different distribution than the one we use to compute the loss. Previous work [29] showed that the Trajectory Balance loss is amenable to off-policy training without importance sampling and weighting without introducing high variance. This is important since importance weighting would require us to weight by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})/q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$, reintroducing the need to compute $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$.
The off-policy benefits of the Trajectory Balance loss provide a strong argument over more common Reinforcement Learning setups. Off-policy training in Reinforcement Learning usually requires importance weighting to be stable, which is not tractable in our setting.
Appendix B Experimental Details
For all experiments, we used as architecture the Graph Convolutional Network [23], with two layers, a hidden size of 256, batch size of 256, and a sampling size of 256 nodes per layer. We implemented the GCNs in GRAPES via PyTorch Geometric [12]. We train for 50 epochs on Cora, Citeseer, and Reddit; 100 epochs on BlogCat, DBLP, Flickr, ogbn-products, Pubmed, snap-patents, and Yelp; and 150 epochs on ogbn-arxiv and ogbn-proteins. The ogbn-proteins and BlogCat datasets do not contain node features, and instead we learn node embeddings for them of dimension 128 for ogbn-proteins, and 64 for BlogCat.
Our experiments were carried out in a single-node cluster setup. We conducted our experiments on a machine with Nvidia RTX A4000 GPU (16GB GPU memory), Nvidia A100 (40GB GPU memory), and Nvidia RTX A6000 GPU (48GB GPU memory) and each machine had 48 CPUs. In total, we estimate that our experiments took 200 compute days.
B.1 Hyperparameter Tuning
We tune the hyperparameters of GRAPES using a random search strategy with the goal of maximizing the accuracy of the validation dataset. We used Weights and Biases for hyperparameter tuning 333 https://wandb.ai . The best-performing hyperparameters for every dataset can be found in our repository https://github.com/dfdazac/grapes. The following are the hyperparameters that we tuned: the learning rate of the GFlowNet, the learning rate of the classification GCN, and the scaling parameter $\alpha$. We used the log uniform distribution to sample the aforementioned hyperparameters with the values from the following ranges respectively, $[1e-6,1e-2]$, $[1e-6,1e-2]$, and [1e2, 1e6]. We kept the other hyperparameters such as the batch size and hidden dimension of the GCN. We used the Adam optimizer [22] for $\operatorname{GCN}_{\operatorname{C}}$ and $\operatorname{GCN}_{\operatorname{S}}$.
B.2 Baselines
Table 3: Differences in experimental setups in related work, obtained from the original publications and their official implementations. *This indicates the total budget in terms of nodes sampled across all layers.
FastGCN [9] LADIES [51] GraphSAINT [47] AS-GCN [18] GAS [13]
Architecture GCN GCN GCN GCN+attention GCN,GCNII,GAT,GIN,APPNP,PNA
Hidden size (16, 128) 256 (128, 256, 512, 2048) (16,256) (256, 512, 1024, 2048)
Number of layers 2 5 (2, 4, 5) 2 (2, 4, 64)
Batch size (256, 1024) 512 (400, 512, 1000) 256 (1,2,5,40,12,100)
Nodes per layer (100, 400) (5, 64, 512) (4500, 6000, 8000)* (128, 256, 512) No sampling
Training epochs (100, 200, 300) 300 2000 (50,100,300) (300, 400, 500, 1000)
For a fair comparison, we adjusted the implementations of the baselines so that the only difference is the sampling methods and the rest of the training conditions are kept the same. In the following, we explain the details of the modifications to each of the baselines.
For LADIES we used the official implementation, which also contains an implementation of FastGCN. We changed the nonlinear activation function from ELU to ReLU, and we removed any linear layers after the two layers of the GCN, set dropout to zero, and disabled early stopping. We also noticed that the original LADIES implementation divided the target nodes into mini-batches, not from the entire graphs as we do, but into random fragments. This means that LADIES and FastGCN do not see all the target nodes in the training data. We kept this setting unchanged because otherwise it significantly slowed down the training of these two methods.
For GraphSAINT we noticed that the GNN consists of two layers of higher order aggregator, which is a combination of GraphSage-mean [16] and MixHop [2], and a linear classification layer at the end. Moreover, the original implementation of GraphSAINT is only applicable to inductive learning on the graphs, where the train graph only contains the train nodes and is entirely different from the validation and test graphs where only the nodes from the validation set and test are available respectively. We argue that in transductive learning, unlike inductive learning, the motivation to scale to larger graphs is higher since the validation and test nodes are also available during training, and therefore, the processed graph is larger. Finally, to keep all the configurations the same as GRAPES, we used PyTorch Geometric’s function for GraphSAINT Random Walk sampler with depth two, which showed the best performance across the three variations of GraphSAINT. Therefore, we use the same GCN and data loader (transductive) as ours and use GraphSAINT to sample a subgraph of nodes for training. We use the node sampler setting since it is the only setting that allows specifying different sampling budgets, and therefore, can be compared to layer-wise methods with the same sampling budget. We also removed the early stopping and use the same number of epochs as GRAPES. Note that for Figure 3, we need to change the sample size while Random Walk sampler can only sample equal to the batch size per layer. Therefore, for these experiments we used the Node Sampler version of GraphSAINT, which allows us to change the sample size. For instance, for sample size $32$ per layer, we set GraphSAINT’s Node Sampler’s sample size to $256+32+32=320$ to sample a subgraph with 320 nodes. Please refer to our repository for more details about the implementation of GraphSAINT.
For GAS, we used the original implementation. However, we changed the configuration of the GCN to have a two-layer GCN with 256 hidden units. We turned off dropout, batch normalization and residual connections in the GCN. We also removed early stopping for the training.
For AS-GCN we removed the attention mechanism used in the GCN classifier. Their method also uses attention in the sampler, which is separate from the classifier, so we keep it.
Appendix C Dataset statistics
We present the statistics of the datasets used in our experiments in Table 4. The splits that we used for Cora, Citeseer, and Pubmed correspond to the “full” splits, in which the label rate is higher than in the “public” splits. For BlogCat, we take the average accuracy of all the methods across the three available splits provided by [48]. For DBLP and snap-patents we use the average of ten random splits because these two datasets had no predefined splits.
Table 4: Statistics of the datasets used in our experiments. The label rate indicates the percentage of nodes used for training. ogbn-proteins and BlogCat do not contain node features, and instead we learn embeddings for nodes in these datasets.
Dataset Task Nodes Edges Features Classes Label Rate (%)
Cora multi-class 2,708 5,278 1,433 7 44.61
CiteSeer multi-class 3,327 4,552 3,703 6 54.91
PubMed multi-class 19,717 44,324 500 3 92.39
Reddit multi-class 232,965 11,606,919 602 41 65.86
DBLP multi-label 28,702 136,670 300 4 80.00
Flickr multi-class 89,250 449,878 500 7 50.00
snap-patents multi-class 2,923,922 27,945,092 269 5 50.00
Yelp multi-label 716,847 6,977,409 300 100 75.00
ogbn-proteins multi-label 132,534 79,122,504 112 65.30
BlogCat multi-label 10,312 667,966 39 60.00
ogbn-arxiv multi-class 169,343 1,157,799 128 40 53.70
ogbn-products multi-class 2,449,029 61,859,076 100 47 8.03
Appendix D Entropy as Node preference measure
Figures 5 and 6 show the mean and standard deviation of entropy in base two of all the datasets. We calculate the mean entropy for as the following:
$E=\frac{1}{n}\sum_{i=1}^{n}p_{i}\cdot\log_{2}(p_{i})+(1-p_{i})\cdot\log_{2}(1-p_{i})$ (8)
where $n$ is the number of neighbors of the nodes sampled in the previous layer and $p_{i}$ is the probability of inclusion for each node, which is the output of the GFlowNet. As the figures show, for certain small datasets (Cora, Citeseer, Pubmed, Flickr) the mean entropy is: 1) very close to 1, indicating that GRAPES prefers every nodes with the probability close to $0.5$, or 2) close to 0 but also with a low standard deviation, meaning that it equally prefers the majority of the nodes with the probability 1 or 0. On the contrary, for the large datasets (Reddit, Yelp, ogbn-arxiv, ogbn-products) by the end of training, the average entropy is lower than 1, with a standard deviation around $0.3$ indicating that GRAPES learns different preferences over different nodes, some with a probability close to 1, and some close to 0.
Figure 5: Combined entropy plots for Citeseer, Cora, Flickr, ogbn-arxiv, ogbn-products, Pubmed, Reddit and Yelp showcasing the mean entropy. The shaded region indicates the standard deviation of the entropy across nodes. The plots compare 1-hop node sampling against 2-hop node sampling.
Figure 6: Combined entropy plots for BlogCat, snap-patents, DBLP and ogbn-proteins, showcasing the mean entropy across epochs. The shaded region indicates the standard deviation of the entropy across nodes. The plots compare 1-hop node sampling against 2-hop node sampling.
Appendix E GPU Memory usage comparison between GRAPES and GAS
We compared different variants of GRAPES, with different sample sizes (32, 256), with GAS [13], which is a non-sampling method. Figure 2 shows the GPU memory allocation (MB), in a logarithmic scale for GRAPES-32, GRAPES-256 and GAS. The three graph methods exhibit distinct performance characteristics across various datasets. We used the max_memory_allocated function in PyTorch to measure the GPU memory allocation. [^4] https://pytorch.org/docs/stable/generated/torch.cuda.max_memory_allocated.html Since this function measures the maximum memory allocation since the beginning of the program, where the memory measurement is done is not relevant.
Appendix F GFlowNet details
This section provides additional details on GFlowNets in general and the GFlowNet version of GRAPES.
F.1 Generative Flow Networks
Generative Flow Networks (GFlowNets) [6, 7] are generative models that can sample from a very large structured space. GFlowNets construct the structure in multiple generation steps. Compared to Reinforcement Learning approaches, GFlowNets learn to sample in proportion to a given reward function, while in Reinforcement Learning, reward functions are maximized. This feature of GFlowNets encourages sampling diverse sets of high-quality structures, instead of only considering the single best structure. GFlowNets have been used in several applications like molecule design and material science [6, 14, 20], Bayesian structure learning [11], scientific discovery [21], and GNN explainability [26]. Similar to our work, the latter utilizes a GFlowNet to sample subgraphs. However, this method explains a trained GNN and is not used to scale GNN training to large graphs.
F.2 GFlowNet and Trajectory Balance Loss
We first give a brief overview of GFlowNets [6, 7] and the trajectory balance loss [28]. Let $\mathcal{G}_{F}=(\mathcal{S},\mathcal{A},\mathcal{S}_{0},\mathcal{S}_{f},R)$ denote a GFlowNet learning problem. Here, $\mathcal{S}$ is a finite set of states that forms a directed graph with $\mathcal{A}$, a set of directed edges representing actions or transitions between states. $\mathcal{S}_{0}\subset\mathcal{S}$ is the set of initial states, $\mathcal{S}_{f}\subset\mathcal{S}$ is the set of terminating states, [^5] Technically, GFlowNets have unique source and terminal (or ‘sink’) states $s_{s}$ and $s_{f}$. The source state has an edge to all initial states, and all terminating states have an edge to the terminal state. and $R:\mathcal{S}_{f}\rightarrow\mathbb{R}_{+}$ is the reward function defined on terminating states. At time $t$, a particular $a_{t}\in\mathcal{A}$ indicates the action taken to transition from state $s^{(t-1)}$ to $s^{(t)}$. A trajectory $\tau$ is a path through the graph from an initial state $s^{(0)}$ to a terminating state $s^{(n)}\in\mathcal{S}_{f}$: $\tau=(s^{(0)}\to\ldots\to s^{(n)})$. A GFlowNet is a neural network that learns to transition from an initial state $s^{(0)}$ to a terminating state where the reward $R(s^{(n)})$ is given. The goal of the GFlowNet is to ensure that following the forward transition probabilities $P_{F}(s^{(t)}|s^{(t-1)})$ leads to final states $s\in\mathcal{S}_{f}$ with probability in proportion to the reward $R$ [6]. The Trajectory Balance (TB) loss [28] is developed with this goal. For a trajectory $\tau=(s^{(0)}\to\ldots\to s^{(n)})$, the TB loss is:
$\mathcal{L}_{TB}(\tau)=\left(\log\frac{Z(s^{(0)})\prod_{t=1}^{n}P_{F}(s^{(t)}|s^{(t-1)})}{R(s^{(n)})\prod_{t=1}^{n}P_{B}(s^{(t-1)}|s^{(t)})}\right)^{2},$ (9)
where $Z:\mathcal{S}_{0}\rightarrow\mathbb{R}_{+}$ computes the total flow of the network from the starting state $s^{(0)}$ and $P_{F}$ and $P_{B}$ are the forward and the backward transition probabilities between the states, where both can be parameterized by a neural network [28].
F.3 GFlowNet Design: States, Actions, and Reward
Next, we explain our choice of $\mathcal{G}_{F}$, that is, the states, actions, terminating states, and reward function, and the coupling of our GFlowNet with the sampling policy $q$. A state $s^{(l-1)}\in\mathcal{S}$ represents a sequence of sets of nodes $s^{(l-1)}=(\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ sampled so far. An action from $s^{(l-1)}$ represents choosing $k$ nodes without replacement among $\mathcal{N}(K^{(l-1)})$. This forms the set of nodes $\mathcal{V}^{(l)}$ in the next layer. Therefore, in an $L$ layer GCN, we construct a sequence of $L$ sets of nodes to reach a terminating state.
We define the optimal sampling policy as having the lowest classification loss in expectation. Therefore, a set of $k$ nodes with a lower classification loss than another set must have a higher probability. Our goal is to design $\mathcal{G}_{F}$ so that it learns the forward and backward transition probabilities proportional to a given reward. We define the reward function as below:
$R(s^{(L)})=R(\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(L)}):=\exp(-\alpha\cdot\mathcal{L}_{\operatorname{C}}(X,Y,K^{(0)},\ldots,K^{(L)})),$ (10)
where $\mathcal{L}_{\operatorname{C}}$ is the classification loss and $\alpha$ is a scaling parameter, which we explain in Section F.4.
Forward Probability. The forward probability $P_{F}(s^{(l-1)}|s^{(l)})$ in GRAPES is the sampling policy $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ defined in Section 4.1.
Backward Probability. Trajectory balance (Equation 9) also requires defining the probability of transitioning backwards through the states. The backward probability is a distribution over all parents of a state. This distribution is not required in our setup, as the state representation $s^{(l)}=(\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l)})$ saves the trajectory taken through $\mathcal{G}_{F}$ to get to $s^{(l)}$. This means the graph for the GFlowNet learning problem $\mathcal{G}_{F}$ is a tree, as each state $s^{(l)}=(\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},\mathcal{V}^{(l)})$ has exactly 1 parent, namely $s^{(l-1)}=(\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$. Since each state has a single parent, we find that $P_{B}(s^{(l-1)}|s^{(l)})=1$ when we retrace the trajectory. We pass the information on when a node is added to the GFlowNet by adding an identifier to the nodes’ embeddings that indicates in what layer it was sampled, as explained in Section 4.1.
Loss derivation Combining our setup with the trajectory balance loss (Equation 9), the GRAPES loss is
$\mathcal{L}_{\operatorname{GFN}}(X,Y,\mathcal{V}^{(0)})=\Big{(}$ $\log\frac{Z(\mathcal{V}^{(0)})\prod_{l=1}^{L}P_{F}(s^{(l)}|s^{(l-1)})}{R(s^{(l)})}\Big{)}^{2}$ (11)
[…]
$\alpha\cdot\mathcal{L}_{\operatorname{C}}(X,Y,K^{(0)},\ldots,K^{(L)})\Big{)}^{2}$
We model the initial-state-dependent normalizer $Z(s^{(0)})$ in Equation 9 with a trainable GCN, namely $\operatorname{GCN}_{Z}(\mathcal{V}^{(0)})$. It predicts the normalizer conditioned on the target nodes given. It is trained together with $\operatorname{GCN}_{\operatorname{S}}$ by minimizing the trajectory balance loss.
We note that, like in GRAPES-RL, we use off-policy sampling from $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$ to train the GFlowNet. See Appendix A for additional details.
F.4 Reward scaling
In our experiments, we noticed that with the bigger datasets, the GFlowNet is more affected by the log-probabilities than the reward from the classification $\operatorname{GCN}_{\operatorname{C}}$. The reason that the term $\log q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ dominates $\mathcal{L}_{\operatorname{GFN}}$ is that $\log q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})=\sum_{i\in\mathcal{N}(K^{(l-1)})}\log p_{i}$, which sums over $|\mathcal{N}(K^{(l-1)})|$ elements. Given a batch size of $256$, this neighborhood can be as big as $52\hskip 1.49994pt000$ nodes, resulting in summing $52\hskip 1.49994pt000$ log-probabilities. The majority of the probabilities $p_{i}$ are values close to zero. Therefore, the above sum would be a large negative number. Since the loss $\mathcal{L}_{\operatorname{C}}$ is, in our experiments, often quite close to zero, the log-probability and its variance dominate the loss. Therefore, we add the hyper-parameter $\alpha$ to the reward and tune it in our experiments.